OpenStudy (anonymous):

Explain, in complete sentences, the difference between a radical equation and an equation simply containing a radical expression. Provide an example of each.

OpenStudy (anonymous):

A radical equation has a variable inside a radical. Equations that have radicals, but not over variables are not radical equations. e.g \[2x=\sqrt{5}\]

OpenStudy (anonymous):

I still do not get it....

OpenStudy (jtvatsim):

Here is an example of a radical equation \[\sqrt{x} = 2\] notice that the x (or variable) is inside the square root. Here is an example of an equation containing a radical expression \[x = \sqrt{2}\] notice that there is a square root here, but the x is not inside it.

OpenStudy (anonymous):

okay thanks guys :)